what are the next three numbers in the sequence: 32,16,8,4,2?? thanks, Camille
Answer by checkley79(3341)

32,16,8,4,2, divide the first 2 numbers 32/16=2
2 divide the next 2 numbers 16/8=2
2 divide the next 2 numbers 8/4=2
2 divide the next 2 numbers 4/2=2
2 thus the multiplier is 1/2
2*1/2=1 for the first missing number
1*1/2=1/2 for the second missing number
1/2*1/2=1/4 for the thirs missing number.
ans= 16,8,4,2,1,1/2,1/4